Sprout is attempting to breed dittany (Origanum dictamnus) with greater oil content in the leaves and flowers, as it is the oil that works to heal wounds. When she measures oil content in the leaves and flowers collected from six different plants, she obtains the following values: 
0.02, 0.03, 0.03, 0.04, 0.06, 0.08 (ml/ g).

Calculate the variance for oil content in dittany. Show all work!
Answer #1

S2= variance

x= term in the data set

n= sample sizeX= Ex ....- (sample mean) = 0.02 +0.03+0.03+0.04 +0.06+0.08. 7 = 0·0633 Sample variance (52) S2 = {(x-x)2 n -1 1 x x x (x-2)

the sample variance for the oil content in dittany is 0.00051 mg/ml.